A triangle has angle measures of 80 degrees and 45 degrees. What would be the measure of the third angle of the triangle
Answer by stanbon(75887)

A triangle has angle measures of 80 degrees and 45 degrees. What would be the measure of the third angle of the triangle
------
3rd = 180-(80+45) = 180-125 = 55 degrees
------------
Cheers,
Stan H.